OpenStudy (anonymous):

The steps to derive the quadratic formula are shown below: Pictures Below Which of the following statements best describes the first incorrect steps: Pictures Below

OpenStudy (anonymous):

steps

OpenStudy (anonymous):

Answers

zepdrix (zepdrix):

Hmm this is weird. I don't see any glaring mistakes. The only one I see, and it's really subtle, step 6 should be:\[\Large\rm x+\frac{b}{2a}=\pm\sqrt{-\frac{c}{a}+\left(\frac{b}{2a}\right)^2}\] Oh oh oh nevermind, it's step 5, they didn't write the perfect square correctly.